Zelnor, Inc., is an all-equity firm with 140 million shares outstanding currently trading for $10.43 per share. Suppose Zelnor decides to grant a total of 14 million new shares to employees as part of a new compensation plan. The firm argues that this new

Zelnor, Inc., is an all-equity firm with million shares outstanding currently trading for per share. Suppose Zelnor decides to grant a total of million new shares to employees as part of a new compensation plan. The firm argues that this new compensation plan will motivate employees and is better than giving salary bonuses because it will not cost the firm anything. Assume perfect capital markets.

a. If the new compensation plan has no effect on the value of Zelnor's assets, what will be the share price of the stock  once this plan is implemented?

b. What is the cost of this plan for Zelnor investors? Why is issuing equity costly in this case?
Answer #1

20210714_143513-min.jpg

9b) step 1) 10.43- 9.48= 0.95           or 14* 9.48= 132.72 or 133

step 2) 140* 0.95= answer: 133


Why is issuing equity costly in this case?

Answer: It's costly because the shareholder equity is being given away to employees for free.






answered by: Andrew San Andres